What is the possibility of drawing either the ace of hearts or the ten of diamonds from a standard deck?
Answer by stanbon(75887)

What is the possibility of drawing either the ace of hearts or the ten of diamonds from a standard deck?
-------------------------
P(aceH or 10D) = P(aceH)+P(10D)= 1/52 + 1/52 = 1/26
----------------------------
Cheers,
Stan H.